OpenStudy (anonymous):

Integral of 33^3(sqrt(x))

OpenStudy (adunb8):

can you rewrite the problem using the equation button on the bottom?

OpenStudy (anonymous):

is this your problem: \(\large \int33^{3\sqrt{x}}dx \) ???

OpenStudy (anonymous):

or this: \(\large \int 33^3 \sqrt{x}dx \) ??

OpenStudy (adunb8):

there are actually the steps there see if you get it sorry if i cant be much of a help =(

OpenStudy (anonymous):

Use wolfram, if you cannot find it by manual means.